# TableDART: Dynamic Adaptive Multi-Modal Routing for Table Understanding

### v1

@@ 4.3 Efficiency Analysis
Inference Efficiency. To quantify the practical benefits of our dynamic adaptive routing policy, we conducted a rigorous efficiency benchmark comparing TableDART’s optimal configuration ($\lambda=0.15$) against a Non-Adaptive Fusion baseline, where every instance is processed through the full multimodal pipeline. The detailed evaluation protocol is in Appendix A.6. The results, presented in Table 3, demonstrate substantial efficiency gains. For example, TableDART averagely achieves a 24.5% reduction in latency compared to the Non-Adaptive Fusion baseline, decreasing the mean inference time from 2.92s to 2.20s per sample. These improvements are a direct consequence of TableDART’s learned routing policy, which intelligently avoids the overhead of the costly Fusion path when a more efficient unimodal path is sufficient.
@@ 4.5 Ablation Study on Key Components
Impact of Dynamic Gating. We ablate the routing strategy to validate our dynamic approach by comparing it with two non-adaptive baselines in Table 4. The results first confirm the effectiveness of our dynamic strategy, as TableDART significantly outperforms the Random Routing baseline. Compared with the Non-Adaptive Fusion baseline, that is, a brute-force strategy where every instance is routed through our LLM Agent-based Fusion path, our method not only achieves comparable results on most benchmarks but also delivers clear gains on datasets such as TABMWP and HiTab. These observations demonstrate the balance of resource-efficiency and performance-effectiveness of our dynamic routing design.
@@
TableDART (Dynamic Routing) 70.58 84.54 62.05 74.37 36.11 81.37 73.59
@@
Impact of Resource-Aware Objective ($\lambda$). We analyze the effect of the resource loss weight $\lambda$, highlighting its two key roles in shaping the final routing policy. (1) Effective Routing Policy Control. As shown in Figure 3, $\lambda$ directly regulates the framework’s behavior. For instance, reducing its value increases reliance on the computationally expensive Fusion path, causing more queries to be routed to the Fusion inference path. (2) Regularization for Improved Generalization. Table 4.5 shows that our proposed Resource-Aware objective also serves as an effective regularizer. The highest performance, both in terms of average results and across four of the seven benchmarks, is not attained with a purely performance-driven policy ($\lambda=0.00$), but rather at a non-zero $\lambda$. By penalizing computational cost, the objective discourages the gate from over-relying on the Fusion path, a tendency that risks overfitting to training artifacts. Instead, it promotes a more balanced routing policy that generalizes more effectively. Based on this observation, we select $\lambda=0.15$ for the final model, which delivers the best average performance at a well-regularized point.
@@ A.1 Reproducibility Statement
We are committed to ensuring the reproducibility of our work. All source code, trained gating network checkpoints, and experiment configurations are provided at the following anonymous GitHub repository: https://anonymous.4open.science/r/TableDART-C52B. Below is a summary of the key components for reproduction.
  • Model Architecture: The overall architecture of TableDART is detailed in Section 3. Specific details of the embedding extraction pipeline and the MLP gating network are provided in Appendix A.3. The constituent models used are TableGPT2-7B (Su et al., 2024), Ovis2-8B (Lu et al., 2024), and Google’s Gemini 2.0 Flash model (Comanici et al., 2025).
@@
  • Datasets and Preprocessing: The datasets used for training and evaluation are described in Section 4. Our training mixture was constructed from five datasets: WTQ (Pasupat & Liang, 2015), TABMWP (Lu et al., 2023), TAT-QA (Zhu et al., 2021), TabFact (Chen et al., 2020), and InfoTabs (Gupta et al., 2020). We evaluate on these five datasets plus two zero-shot datasets: HiTab (Cheng et al., 2022) and FeTaQA (Nan et al., 2022). Our specific data construction protocol is detailed in Section 4.1 and Appendix A.5.
  • Hyperparameters and Computational Infrastructure: A comprehensive list of all hyperparameters is provided in Table 10. All experiments were conducted on a single NVIDIA H100 80GB GPU. Full details on the training pipeline, hyperparameters, and computational setup are located in Appendix A.5.
@@ A.3.1 Embedding Architecture and Computational Analysis
  • Question Embedding: 384 dimensions from all-MiniLM-L6-v2 sentence transformer. This pre-trained model provides question-type agnostic semantic representations that complement the features from the single-modality models, enabling the gating network to learn routing patterns based on question semantics rather than model-specific encodings.
@@ A.3.2 Gating Network Architecture Specification
The MLP gating network implements a standard two-layer architecture designed for efficient and effective routing. The detailed forward pass is defined as:
@@
Implementation Details. The hidden dimension is set to 256, a choice empirically found to balance representational capacity for learning complex routing patterns with computational efficiency. Dropout with a probability of $p=0.1$ is applied after the hidden layer to prevent overfitting. We use the ReLU activation function for its computational efficiency and stable gradient properties. All weight matrices are initialized using Xavier uniform initialization (Glorot & Bengio, 2010), and bias vectors are initialized to zero to ensure stable training dynamics.
@@ A.5.1 Datasets and Metrics
Table Question
Answering (TQA)
TABMWP 2,000 300 7,686
@@
Table Fact
Verification (TFV)
TabFact 2,000 300 6,845
@@ A.5.2 Comparative Models and Rationale
  • Constituent Models: To demonstrate the necessity of a dynamic routing framework, we evaluate our constituent single-modality models in isolation: the Table-as-Text model, TableGPT2-7B (Su et al., 2024), and the Table-as-Image model, Ovis2-8B (Lu et al., 2024). This establishes a performance baseline and highlights scenarios where each modality excels individually.
@@
  • Broader Competitive Landscape: To position TableDART within the broader field, we benchmark it against an extensive suite of single-modality baselines in the 7-8B parameter range. Our selection of Table-as-Text models includes generalist LLMs like Llama-2-7B (Touvron et al., 2023) and Llama3-Instruct-8B Dubey et al. (2024), as well as the specialized TableLlama-7B (Zhang et al., 2024). The set of Table-as-Image MLLMs is equally comprehensive, featuring MiniGPT-4-7B (Zhu et al., 2024a), mPLUG-Owl-7B (Ye et al., 2023), mPLUG-Owl2-7B (Ye et al., 2024), LLaVA v1.5-7B (Liu et al., 2024), Table-LLaVA-7B (Zheng et al., 2024), Qwen-VL-7B Bai et al. (2023), InternLM-XComposer2-7B (Zhang et al., 2023), Monkey-7B (Li et al., 2024), and MiniCPM-V-2.6-8B (Yao et al., 2024).
@@ A.5.3 Training and Implementation Details
Our training set is a 10,000-sample mixture, constructed by randomly sampling 2,000 instances from five datasets: TABMWP, WTQ, TAT-QA, TabFact, and InfoTabs (Lu et al., 2023; Pasupat & Liang, 2015; Zhu et al., 2021; Chen et al., 2020; Gupta et al., 2020). Following the protocol of HIPPO, FeTaQA and HiTab are excluded from training due to their non-accuracy-based evaluation metric (BLEU) and challenges in serializing complex hierarchical structures, respectively (Nan et al., 2022; Cheng et al., 2022). An additional 15% validation split (1,500 samples) is generated using the same random sampling procedure with a fixed seed to ensure reproducibility.
@@
The resource cost vector $\mathbf{c}$ used in our resource-aware objective is derived from empirical measurements to provide a realistic estimate of each path’s computational expense. We measured these costs on a representative testbed of 70 samples (10 from each benchmark), averaging over 10 timed runs following 5 warm-up iterations. Our final cost metric offers a holistic balance of latency (seconds per instance) and throughput (tokens per second), defined as: $\text{Cost}=0.5\times(\text{Avg Latency})+0.5\times(1.0/\text{Avg TPS})$. For the Fusion path, its latency assumes parallel execution of its base models plus a 0.3s API average overhead. The resulting empirically derived cost values are summarized in Table 9.

### v2

@@ 4.3 Efficiency Analysis
Inference Efficiency. To quantify the practical benefits of dynamic adaptive routing, we benchmark TableDART’s optimal configuration ($\lambda=0.15$) against a Non-Adaptive Fusion baseline that processes every instance via the full multimodal pipeline. Both settings use identical backbone models and Fusion agents, ensuring a fair comparison to isolate the efficiency gains of our adaptive policy. The detailed evaluation protocol is in Appendix A.6. The results, presented in Table 3, demonstrate substantial efficiency gains. For example, TableDART achieves an average reduction of 24.5% in latency compared to the Non-Adaptive Fusion baseline, decreasing the mean inference time from 2.92s to 2.20s per sample. These improvements are a direct consequence of TableDART’s learned routing policy, which reserves the computationally expensive Fusion path for instances that truly require multimodal reasoning, while assigning simpler cases to the more efficient unimodal paths. As shown in Appendix A.7 and Appendix B.3, the model routes multimodally complex instances to the Fusion path while assigning simpler ones to the appropriate unimodal path. In particular, Figure 6 shows that 34.8% of TAT-QA instances cannot be solved by either single-modality model alone, and Figure 10 confirms that the learned policy routes 88.7% of these instances to the powerful Fusion path. In contrast, for simpler datasets such as TABMWP, where single-modality models are highly effective, the routing policy assigns 97.2% of the instances to the more efficient Image-only path. Overall, TableDART’s inference efficiency derives from its intelligent dynamic routing mechanism, which preserves effectiveness while substantially reducing unnecessary computational cost.
@@ 4.5 Ablation Study on Key Components
Impact of Dynamic Gating. We ablate the routing strategy to validate our dynamic approach by comparing it with two non-adaptive baselines in Table 4. The results first confirm the effectiveness of our dynamic strategy, as TableDART significantly outperforms the Random Routing baseline. Compared with the Non-Adaptive Fusion baseline, which uses the same backbone models as TableDART but applies a brute-force strategy where every instance is routed through our LLM Agent-based Fusion path, our method not only achieves comparable results on most benchmarks but also delivers clear gains on datasets such as TABMWP and HiTab. This indicates that TableDART’s effectiveness is not solely attributable to the strong Fusion agent, but arises from the dynamic routing mechanism itself. As shown in Figure 10, the learned policy routes most instances to the appropriate unimodal paths and uses Fusion only when needed. In many cases, a single modality already provides sufficient information, and forcing unnecessary multimodal processing introduces extra cost and potential noise, which is why TableDART surpasses the Non-Adaptive Fusion baseline on several datasets. Overall, these observations demonstrate the balance of resource-efficiency and performance-effectiveness of our dynamic routing design.
@@
TableDART (Dynamic Routing) 70.58 84.54 62.05 74.37 36.11 81.37 76.22
@@
Impact of Resource-Aware Objective ($\lambda$). We analyze the effect of the resource loss weight $\lambda$, highlighting its two key roles in shaping the final routing policy. (1) Effective Routing Policy Control. As shown in Figure 3, $\lambda$ directly regulates the framework’s behavior. For instance, reducing its value increases reliance on the computationally expensive Fusion path, causing more queries to be routed to the Fusion inference path.
(2) Regularization for Improved Generalization. Table 4.5 shows that our proposed resource-aware objective also serves as an effective regularizer. The highest performance, both in terms of average results and across four of the seven benchmarks, is not attained with a purely performance-driven policy ($\lambda=0.00$), but rather at a non-zero $\lambda$. By penalizing computational cost, the objective discourages the gate from over-relying on the Fusion path, a tendency that risks overfitting to training artifacts. Instead, it promotes a more balanced routing policy that generalizes more effectively. In addition, varying $\lambda$ has a direct impact on inference efficiency. As illustrated in Appendix B.4 (Figure 11 and Table 11), efficiency generally improves as $\lambda$ increases, since a larger penalty discourages unnecessary use of the Fusion path. However, the trend is not strictly monotonic, because each $\lambda$ induces a distinct routing policy that balances accuracy and computational cost differently. Based on this observation, we select $\lambda=0.15$ for the final model, which achieves the second-best average accuracy (74.86%, within 0.19 points of the best) while requiring 8.4% less inference latency, representing a well-regularized operating point on the performance–efficiency frontier. This configuration enables a truly adaptive policy, which tailors routing to the task type, directing 97.2% of queries to the Image-only path for the structure-heavy TABMWP, while shifting to 67.5% Text-only for the semantics-focused InfoTabs (see Appendix B.3 and Fig. 10 for a full breakdown).
@@ 5 Conclusion
Acknowledgments
The Australian Research Council partially supports this work under the streams of Future Fellowship (Grant No. FT210100624), the Discovery Project (Grant No. DP240101108 and DP260100326), and the Linkage Project (Grant No. LP230200892 and LP240200546).
@@ A.1 Reproducibility Statement
We are committed to ensuring the reproducibility of our work. All source code, trained gating network checkpoints, and experiment configurations are provided at the following GitHub repository: https://github.com/xiaobo-xing/TableDART. Below is a summary of the key components for reproduction.
  • Model Architecture: The overall architecture of TableDART is detailed in Section 3. Specific details of the embedding extraction pipeline and the MLP gating network are provided in Appendix A.3. The primary constituent models are TableGPT2-7B (Su et al., 2024) and Ovis2-8B (Lu et al., 2024), with Qwen2.5-VL-7B (Bai et al., 2025) additionally employed to validate generalization. Google’s Gemini 2.0 Flash model (Comanici et al., 2025) implements the Fusion agent.
@@
  • Datasets and Preprocessing: The datasets used for training and evaluation are described in Section 4. Our training mixture was constructed from five datasets: WTQ (Pasupat and Liang, 2015), TABMWP (Lu et al., 2023), TAT-QA (Zhu et al., 2021), TabFact (Chen et al., 2020), and InfoTabs (Gupta et al., 2020). We evaluate on these five datasets plus two zero-shot datasets: HiTab (Cheng et al., 2022) and FeTaQA (Nan et al., 2022). Our specific data construction protocol is detailed in Section 4.1 and Appendix A.5.
  • Hyperparameters and Computational Infrastructure: A comprehensive list of all hyperparameters is provided in Table 10. All experiments were conducted on a single NVIDIA H100 80GB GPU. These computational resources were provided by the Bunya supercomputer (The University of Queensland Research Computing Centre, 2024), operated by The University of Queensland Research Computing Centre (RCC). Full details on the training pipeline, hyperparameters, and computational setup are located in Appendix A.5.
@@ A.3.1 Embedding Architecture and Computational Analysis
  • Question Embedding: 384 dimensions from all-MiniLM-L6-v2 (Reimers and Gurevych, 2019) sentence transformer. This pre-trained model provides question-type agnostic semantic representations that complement the features from the single-modality models, enabling the gating network to learn routing patterns based on question semantics rather than model-specific encodings.
@@ A.3.2 Gating Network Architecture Specification
Following standard practice in dynamic routing and mixture-of-experts systems in the prior works (Lepikhin et al., 2021; Zhu et al., 2022; Xue and Marculescu, 2023), the MLP gating network implements a standard two-layer architecture designed for efficient and effective routing. The detailed forward pass is defined as:
@@
Implementation Details. The hidden dimension is set to 256, a choice empirically found to balance representational capacity for learning complex routing patterns with computational efficiency. Dropout with a probability of $p=0.1$ is applied after the hidden layer to prevent overfitting. We use the ReLU activation function for its computational efficiency and stable gradient properties. All weight matrices are initialized using Xavier uniform initialization (Glorot and Bengio, 2010), and bias vectors are initialized to zero to ensure stable training dynamics.
@@ A.5.1 Datasets and Metrics
Table Question Answering (TQA) TABMWP 2,000 300 7,686
@@
Table Fact Verification (TFV) TabFact 2,000 300 6,845
@@ A.5.2 Comparative Models and Rationale
  • Constituent Models: To demonstrate that TableDART’s performance stems from its dynamic framework, we evaluate its core constituent models as standalone baselines. These include: (1) the Table-as-Text model, TableGPT2-7B (Su et al., 2024); (2) the primary Table-as-Image model, Ovis2-8B (Lu et al., 2024); (3) the additional visual expert used for generalization, Qwen2.5-VL-7B (Bai et al., 2025); and (4) the backbone of our Fusion agent, Google Gemini 2.0 Flash (Comanici et al., 2025). We selected these backbone models because they are recent and strong representatives of their respective paradigms. Evaluating the single-modality experts in isolation establishes a performance baseline. Specifically, including Qwen2.5-VL-7B enables us to test the framework’s generalization capabilities across different visual backbones, while comparing against Gemini as a standalone MLLM verifies that our performance lift arises from the routing mechanism rather than only the capacity of the Fusion agent alone.
@@
  • Broader Competitive Landscape: To position TableDART within the broader field, we benchmark it against an extensive suite of single-modality baselines in the 7-8B parameter range. Our selection of Table-as-Text models includes generalist LLMs like Llama-2-7B (Touvron et al., 2023) and Llama3-Instruct-8B (Dubey et al., 2024), as well as the specialized TableLlama-7B (Zhang et al., 2024). The set of Table-as-Image MLLMs is equally comprehensive, featuring MiniGPT-4-7B (Zhu et al., 2024a), mPLUG-Owl-7B (Ye et al., 2023), mPLUG-Owl2-7B (Ye et al., 2024), LLaVA v1.5-7B (Liu et al., 2024), Table-LLaVA-7B (Zheng et al., 2024), Qwen-VL-7B (Bai et al., 2023), InternLM-XComposer2-7B (Zhang et al., 2023), Monkey-7B (Li et al., 2024), TabPedia-7B (Zhao et al., 2024), SynTab-LLaVA-7B (Zhou et al., 2025), Qwen2.5-VL-7B (Bai et al., 2025) and MiniCPM-V-2.6-8B (Yao et al., 2024).
For a comprehensive and fair comparison, we report several baseline results directly from their original publications. Specifically, the results for the following models are adopted from the HIPPO paper (Liu et al., 2025): Llama-2-7B, Llama3-Instruct-8B, TableLlama-7B, MiniGPT-4-7B, mPLUG-Owl-7B, mPLUG-Owl2-7B, LLaVA v1.5, Table-LLaVA-7B, MiniCPM-V-2.6-8B, Qwen-VL-7B, InternLM-XComposer2-7B, Monkey-7B, and the HIPPO model itself. Furthermore, the results for TableGPT2-7B are sourced from its proposal paper (Su et al., 2024), and the results for SynTab-LLaVA are sourced from Zhou et al. (2025). We also adopt the results for TabPedia and Ovis2-8B (except TABMWP dataset) from Jiang et al. (2025). All other model results not mentioned above were generated by our own experimental runs.
@@ A.5.3 Training and Implementation Details
Our training set is a 10,000-sample mixture, constructed by randomly sampling 2,000 instances from five datasets: TABMWP, WTQ, TAT-QA, TabFact, and InfoTabs (Lu et al., 2023; Pasupat and Liang, 2015; Zhu et al., 2021; Chen et al., 2020; Gupta et al., 2020). Following the protocol of HIPPO (Liu et al., 2025), FeTaQA and HiTab are excluded from training due to their non-accuracy-based evaluation metric (BLEU) and challenges in serializing complex hierarchical structures, respectively (Nan et al., 2022; Cheng et al., 2022). An additional 15% validation split (1,500 samples) is generated using the same random sampling procedure with a fixed seed to ensure reproducibility.
@@
The resource cost vector $\mathbf{c}$ used in our resource-aware objective is derived from empirical measurements to provide a realistic estimate of each path’s computational expense. We measured these costs on a representative testbed of 70 samples (10 from each benchmark), averaging over 10 timed runs following 5 warm-up iterations. Our final cost metric offers a holistic balance of latency (seconds per instance) and throughput (tokens per second), defined as: $\text{Cost}=0.5\times(\text{Avg Latency})+0.5\times(1.0/\text{Avg TPS})$. For the Fusion path, its latency assumes parallel execution of its base models plus a 0.3s API average overhead, reflecting a realistic deployment scenario. The resulting empirically derived cost values are summarized in Table 9.
@@ B.2 Analysis of the Learned Routing Policy
B.3 Analysis of Per-Dataset Routing Policies
To provide quantitative evidence of TableDART’s adaptive capabilities, we analyze the routing decisions of our final model (trained with $\lambda=0.15$) on each of the seven test benchmarks. Figure 10 visualizes the learned policy, revealing how the framework dynamically allocates resources based on the specific demands of each dataset.
The analysis highlights several key adaptive behaviors:
  • Adaptation to Modality Strengths: The model demonstrates a profound understanding of modality-task alignment. For TABMWP, a benchmark requiring structural and mathematical reasoning, the policy routes an overwhelming 97.2% of instances to the Image-only path. Conversely, for InfoTabs, a fact-verification task dependent on fine-grained semantics, the strategy shifts dramatically to favor the Text-only path (67.5%).
  • Adaptation to Task Difficulty: On challenging benchmarks where single-modality models are likely to fail, such as TAT-QA, the policy learns to prioritize correctness by invoking the powerful but costly Fusion path for 88.7% of cases. This demonstrates an ability to gauge task difficulty and escalate to a more robust strategy when necessary.
  • Balanced, Nuanced Strategies: For benchmarks with a mix of challenges like WTQ and FeTaQA, the model learns a more balanced policy, distributing queries across all three paths. This indicates that the routing is not coarse-grained but makes fine-tuned, instance-level decisions.
In summary, this analysis provides compelling evidence that TableDART operates as a truly adaptive framework, which is the key to its state-of-the-art performance and efficiency.
Figure 10: The learned routing policy distribution for each benchmark using the final TableDART model ($\lambda=0.15$). The distinct strategies across datasets provide strong evidence of the framework’s adaptive nature.
B.4 Effect of $\lambda$ on Inference Efficiency
To more clearly characterize how the resource-aware objective influences efficiency, we report empirical latency and throughput measurements across different values of $\lambda$ in Table 11. As shown, $\lambda$ has a direct impact on inference cost. While larger $\lambda$ generally encourages more efficient routing by discouraging unnecessary use of the Fusion path, the trend is not strictly monotonic. This is because each $\lambda$ induces a distinct routing policy that balances accuracy and computational cost differently.
Figure 11 visualizes this relationship by plotting average latency against average accuracy. The resulting curve illustrates a natural performance–efficiency frontier: smaller $\lambda$ values (e.g., $0.0$ and $0.05$) tend to overuse the Fusion path, while very large values (e.g., $1.0$) favor efficiency at the cost of reduced accuracy. The configuration $\lambda=0.15$ provides the best overall balance, achieving the second-highest average accuracy (within 0.19 points of the best) while maintaining strong efficiency (8.4% less inference latency).
Figure 11: Performance–efficiency trade-off across different values of the resource loss weight ($\lambda$). The red star marks the selected configuration.
Table 11: Inference efficiency for different values of $\lambda$. Higher values encourage more efficient routing, though the effect is not strictly monotonic due to the distinct routing policies induced by each setting.
$\lambda$ Latency (s) Throughput (TPS)
0.00 $2.3675$ $15.14$
0.05 $2.4021$ $13.72$
0.10 $2.0832$ $23.85$
0.15 $2.2008$ $17.77$
1.00 $1.8490$ $29.56$